Manhattan Partners provides management consulting services to government and corporate clients. Concord has two support departments---administrative services​ (AS) and information system ​(IS)---and two operating departments---government consulting​ (GOVT) and corporate consulting​ (CORP). For the first quarter of 2017, Manhattan's cost records indicate the​ following:

SUPPORT

OPERATING

AS

IS

GOVT

CORP

Total

Budgeted overhead costs before any

interdepartment cost allocations

$390,000

$3,600,000

$8,750,000

$12,540,000

$25,280,000

Support work supplied by AS

(budgeted head count)

---

20%

48%

32%

100%

Support work supplied by IS

(budgeted computer time)

10%

---

36%

54%

100%

The two support​ departments' costs have been allocated to the two operating departments using the direct method and the​ step-down method​ (AS department​ first) and the​ step-down method​ (IS department​ first). The results of the allocations are as​ follows:

direct method

Support Departments

Operating Departments

Direct Method

AS

IS

GOVT

CORP

Total

Budgeted overhead costs

before interdepartment cost allocations

$390,000

$3,600,000

$8,750,000

$12,540,000

$25,280,000

Allocation of AS costs

(390,000)

234,000

156,000

Allocation of IS costs

(3,600,000)

1,440,000

2,160,000

Total budgeted overhead of operating departments

$0

$0

$10,424,000

$14,856,000

$25,280,000

step down method (AS) first

Support Departments

Operating Departments

Step-down Method

AS

IS

GOVT

CORP

Total

Budgeted overhead costs

before interdepartment cost allocations

$390,000

$3,600,000

$8,750,000

$12,540,000

$25,280,000

Allocation of AS costs

(390,000)

78,000

187,200

124,800

$0

3,678,000

Allocation of IS costs

(3,678,000)

1,471,200

2,206,800

Total budgeted overhead of operating departments

$0

$10,408,400

$14,871,600

$25,280,000

step down method (IS) first

Support Departments

Operating Departments

Step-down Method

AS

IS

GOVT

CORP

Total

Budgeted overhead costs

before interdepartment cost allocations

$390,000

$3,600,000

$8,750,000

$12,540,000

$25,280,000

Allocation of IS costs

360,000

(3,600,000)

1,296,000

1,944,000

$750,000

$0

Allocation of AS costs

(750,000)

450,000

300,000

Total budgeted overhead of operating departments

$0

$10,496,000

$14,784,000

$25,280,000

summary of the three methods

GOVT

CORP

Direct method

$1,674,000

$2,316,000

Step-down (AS first)

1,658,400

2,331,600

Step-down (IS first)

1,746,000

2,244,000

Requirement 1a. Allocate the two support​ departments' costs to the two operating departments using the reciprocal method.

​First, determine the linear equations for each support department.

___

Now allocate the support​ departments' costs​ (AS and IS department costs that you determined in the prior​ step) to the two operating departments using the reciprocal method. ​(Round decimal calculations to three decimal​ places, .XXX, and your final answers to the nearest whole dollar. Use parentheses or a minus sign when decreasing departments by allocating costs. Enter a​ "0" for any zero​ balances.)

Requirement 1b. Allocate the two support​ departments' costs to the two operating departments using the reciprocal method. Use repeated iterations. ​(Round your final answers to the nearest whole dollar. Use parentheses or a minus sign when decreasing departments by allocating costs. Enter a​ "0" for any zero​ balances.)
Answer #1

SUPPOSE AS COST =X

IS COST = Y

X = 390000+ 0.10Y   EQUATION 1

Y = 3600000+0.20X    EQUATION 2

Y = 3600000+0.20(390000+0.10Y)

Y = 3600000+78000+0.02Y

0.98Y = 3678000

Y =3753061

X = 390000+0.10(3753061)

= 765306

RECIPROCAL METHOD
AS IS GOVT CORP TOTAL
INTERDEPARTMENTAL COST ALLOCATION 390000 3600000 8750000 12540000 25280000
AS COST ALLOCATED -765306

153061

[765306*20%]

367347

[765306*48%]

244898

[765306*32%]

IS COST ALLOCATED

375306

[3753061*10%]

-3753061

1351102

[3753061*36%]

2026653

[3753061*54%]

TOTAL 0 0 10468449 14811551 25280000